OpenStudy (anonymous):

perform indicated operations. simplify all answers (2y^2-7y-15/42y^2-29y-5) x (12y^2-16y+5/7y^2-36y+5) / (4y^2-9/49y^2-1) HELP PLEASE!

OpenStudy (anonymous):

2\[\frac{ 2y^{2}-7y-15}{ 42y^{2}-29y-5}\times \frac{ 12y ^{2}-16y+5 }{7y ^{2}-36y+5}\div \frac{ 4y ^{2}-9 }{49y ^{2}-1 }\]

OpenStudy (owlcoffee):

First, perform the multiplication, which is just denominator, multiplied denominator and numerator multiplied numerator.

OpenStudy (anonymous):

i ll try that! thank you :)

OpenStudy (owlcoffee):

After you do that, you can focus on the division.

OpenStudy (loser66):

To me, you should try to factor the last term first (both numerator and denominator), then make reciprocal multiplication. I meant: \[.....\div \dfrac{4y^2-9}{49y^2-1} = ....*\dfrac{49y^2-1}{4y^2-9}\]

OpenStudy (loser66):

Now, factor them, it becomes \(....*\dfrac{(7y+1)(7y-1)}{(2y-3)(2y+3)}\) So that, I am pretty sure that the first 2 terms can have factors such that we can cancel out with the last term. For example, the first term \(\dfrac{2y^2-7y-15}{42y^2-29y-5}=\dfrac{(2y+3)(y-5)}{(7y+1)(6y-5)}\) Do the same with the middle one, then cancel out to get the answer

OpenStudy (anonymous):

thank you that is very helpful.. i was thinking about factoring them as well.

OpenStudy (anonymous):

thank you!

OpenStudy (loser66):

np